At what temperature have the Celsius and Fahrenheit readings the same value


1
Expert's answer
2020-11-09T14:10:55-0500

We have 

Tf = 9/5 Tc +32 


where Tf is temperature in fahrenheit and Tc in degree celsius


when

Tf = Tc


Tf = 9/5 Tf +32 


-4/5 Tf = 32


Tf = -40 F


Tc = -40 C


Answer


At −40 oC

  •